OpenStudy (anonymous):

What is the slope of a line is parallel to y = (1/2)x + 3?

OpenStudy (anonymous):

can you explain? :{

OpenStudy (texaschic101):

In y = mx + b form, the slope of the line is in the m position. Can you tell me the slope of y = 1/2x + 3 ?

OpenStudy (texaschic101):

y = mx + b y = (1/2)x + b now what number is in the m position ?

OpenStudy (texaschic101):

ok...y = 1/2x + 3 the number in the m position is 1/2. That is you slope. A parallel line will have the same slope. So your parallel line slope is also 1/2